TUGBOAT

Volume 16, Number 1 ^{/} March 1995

3 Addresses 4 EuroTEX94 Contest Answers ^{/} Włodek Bzyl

General Delivery 5 Opening words ^{/} Christina Thiele and Michel Goossens 8 Editorial comments ^{/} Barbara Beeton

Software & Tools 9 Making MakeTeXPK safer for Unix installations ^{/} Michael Jaegermann 12 Hyphenation exception log ^{/} Barbara Beeton

Philology 18 Configuring TEX or L^{A}TEX for typesetting in several languages ^{/} Claudio Beccari

30 How to make a foreign language pattern file: Romanian ^{/} Claudio Beccari,

Radu Oprea and Elena Tulei

42 TEX and Linguistics ^{/} Christina Thiele

Font Forum 45 Introducing ^{/} Alan Hoenig 46 Some techniques ^{/} Yannis Haralambous 54 The program a2ac — Font handling on the PostScript level ^{/} Petr Olsak 60 Problems of the conversion of fonts to PostScript Type 1 ^{/}

Basil Malyshev

69 Partial font embedding utilities for PostScript Type-1 fonts ^{/}

Basil Malyshev and Michel Goossens

78 Tight setting with TEX ^{/} Alan Jeffrey

L^{A}TEX 80 XΥMTEX for drawing chemical structural formulas ^{/} Shinsaku Fujita

News & Announcements 89 Calendar 90 Production notes ^{/} Mimi Burbank 91 Coming next issue

TUG Business 92 Institutional members

Forms 93 TUG membership application

Advertisements 94 TEX consulting and production services